# JWST/NIRSpec insights into the circumnuclear region of Arp 220: A detailed kinematic study

### v1

@@
The study of starburst and AGN feedback is crucial for understanding the regulation of star formation and the evolution of galaxies across cosmic time. Arp 220, the closest ultraluminous infrared galaxy (ULIRG), is in an advanced phase of a major merger with two distinct nuclei, and shows evidence of multi-phase (molecular, ionised, neutral) and multi-scale (from ¡ 0.1 to ¿ 5 kpc) outflows. Therefore, it represents an ideal system for investigating outflow mechanisms and feedback phenomena in detail. Using new JWST NIRSpec IFU observations, we investigate the spatially resolved gaseous (in both ionized and hot molecular phases) and stellar kinematics in the innermost 1 kpc. We decouple the different gas kinematic components through multi-Gaussian fitting, identifying distinct multi-phase outflows associated with the two nuclei, with velocities up to km s-1, as well as two counter-rotating discs around each nucleus embedded in a larger-scale rotational disc. We compute the mass ( M⊙), mass outflow rate ( M⊙ yr-1) and energetics ( erg s-1) for each outflow, finding that the ionized and hot molecular outflowing gas contribute around 2-30% to the total mass and the energy of the outflows, as inferred from the combination of multi-wavelength information. We discuss the possible origin of the outflows, finding no compelling evidence to prefer a starburst or AGN driven scenario. Regardless of their nature, outflows in Arp 220 propagate in multiple directions from parsec to kiloparsec scales, potentially impacting a significant portion of the host galaxy. This contrasts with isolated systems where outflows typically follow a more collimated path, and do not affect the interstellar medium throughout the entire galaxy. This study highlights the importance of investigating merging systems with multi-wavelength facilities, including JWST/NIRSpec IFS, to obtain a comprehensive understanding of feedback mechanisms in galaxy evolution.
@@
Ultraluminous infrared galaxies (ULIRGs), with infrared luminosities LIR ¿ L⊙, are among the most luminous objects in the local Universe. Their extreme IR luminosity is due to the combination of high star formation rate (SFR) and active galactic nuclei (AGN) activity, coupled with a high dust content, which intercept the UV photons emitted from either hot young stars or AGN and re-emit them at longer wavelengths (Soifer et al. 1984; Sanders & Mirabel 1996; Lonsdale et al. 2006; U et al. 2012 and references therein).
@@
In the starburst-quasar evolutionary scenario (e.g., Sanders et al. 1988; Hopkins et al. 2008), ULIRGs are thought to represent a rapidly growing phase of massive galaxies. Within this framework, the merger of gas-rich galaxies triggers a phase of vigorous star formation, fueling the growth of supermassive black holes (SMBHs) and the eventual onset of AGN activity. As the merger progresses, powerful outflows generated by both starburst and AGN winds shape the galactic environment, influencing subsequent star formation and black hole accretion with feedback mechanisms (e.g. Di Matteo et al. 2005; Silk 2013).
@@
Galactic outflows are expected to be more important at high redshift (), during the peak of the cosmic star formation and SMBHs accretion (Madau & Dickinson 2014), and hence where the feedback mechanisms may be maximized (e.g. Brusa et al. 2015; Talia et al. 2017; Perrotta et al. 2019; Kakkad et al. 2020; Cresci et al. 2023). Remarkably, recent observations have revealed some similarities between local ULIRGs and distant starbursts at redshift (e.g. Arribas et al. 2012; Hung et al. 2014; Bellocchi et al. 2022; McKinney et al. 2023). They highlight the importance of understanding the outflow phenomena occurring in local ULIRGs: in fact, the proximity of these sources allows for more detailed observations, characterised by angular resolution and signal-to-noise ratios (S/N) higher than those focusing on more distant galaxies.
@@
Several studies have demonstrated the presence of multi-phase (ionised, neutral, and molecular) gas outflows in ULIRGs (e.g. Sturm et al. 2011; Bellocchi et al. 2013; Veilleux et al. 2013; Arribas et al. 2014; Feruglio et al. 2015; Cazzoli et al. 2016; Fluetsch et al. 2021; Perna et al. 2021; Lai et al. 2022; Bohn et al. 2024). In this work we focus on the closest and prototypical ULIRG Arp 220, where evidence of multi-phase outflows has already been obtained from several multi-wavelength observations (e.g. Arribas et al. 2001; McDowell et al. 2003; Colina et al. 2004; Perna et al. 2020; Wheeler et al. 2020; Lamperti et al. 2022; Ueda et al. 2022).
@@
Arp 220 is located at a distance of Mpc, and has a luminosity log () = 12.2 (e.g. Pereira-Santaella et al. 2021), and a stellar mass of log (/ (U et al. 2012). It is a late-stage merger, containing two compact ( pc), highly obscured nuclei (east and west) separated by 370 pc (1″) which dominate the IR luminosity (e.g. Scoville et al. 2007). The two nuclei are sites of intense star formation, with SFR M⊙ yr-1 (e.g. Nardini et al. 2010; Varenius et al. 2016) and extremely high SFR surface densities ( M⊙ yr-1 kpc2, from radio and far-IR observations; e.g. Barcos-Muñoz et al. 2015; Pereira-Santaella et al. 2021). There is still no convincing direct evidence for the presence of AGN activity in either of the two nuclei (see Perna et al. 2024 and references therein).
@@
This paper presents a kinematic examination of the nuclear region of Arp 220, observed with the Integral Field Spectrograph (IFS) unit of the Near Infrared Spectrograph (NIRSpec) instrument on board the James Webb Space Telescope (JWST) (Jakobsen et al. 2022; Böker et al. 2022), collected as part of the JWST/NIRSpec IFS Guaranteed Time Observations (GTO) survey ”Resolved structure and kinematics of the nuclear regions of nearby galaxies” (program lead: Torsten Böker). Our goal is to investigate the spatially resolved gaseous and stellar kinematics in the innermost 1 kpc of Arp 220, and better understand the feedback processes affecting this ULIRG system.
@@
This manuscript is organised as follows. Section 2 presents the NIRSpec observations and data reduction. Sections 3 and 4 present the spectral fitting analysis and the results through the characterization of the spatially resolved kinematics of the atomic and molecular gas and the identification and study of the different structures found in the central 1.5 1.8 kpc of Arp 220. In Sect. 5 we describe the main properties of the ISM. In Sect. 6 we discuss our results, and the comparison with the literature. Finally, Sect. 7 summarises our conclusions. Throughout this work, we assume and km s-1 Mpc-1 (Bennett et al., 2014). With this cosmology, corresponds to 0.368 kpc at the distance of Arp 220.
@@ 2 Observation and data reduction
Arp 220 was observed using the JWST/NIRSpec instrument in IFS mode (Böker et al. 2022), as part of the JWST GTO program 1267 (PI: D. Dicken), on March 6, 2023.
The dataset comprises two distinct pointings, each employing the first four positions of the medium cycling dithering pattern with 15 groups per integration and one integration per exposure, using the NRSIRS2RAPID readout pattern. The first set is centered at the midpoint between the coordinates of the western (W) and eastern (E) nuclei, while the second set is offset by approximately 1″ towards the northwest to encompass the H and [N ii] emission shell (Lockhart et al. 2015; Perna et al. 2020). Each set has a total integration time of 933 seconds and utilises three high-resolution (R ) configurations: G140H/F100LP, G235H/F170LP, and G395H/F290LP, covering wavelength ranges of 0.97 – 1.82 m, 1.66 – 3.05 m, and 2.87 – 5.14 m, respectively. However, this paper focuses solely on the first two configurations, as our objective is to study the kinematics of the emitting gas, with its strongest transitions occurring at wavelengths below 2.4 m.
The data reduction of Arp 220 observations is described in detail in Perna et al. (2024); here we summarise the relevant information. We used the JWST pipeline v1.8.2 with CRDS context 1063. A patch is included to correct some important bugs that affect this specific version of the pipeline; noise and outliers are corrected following the approaches described in Perna et al. (2023) and D’Eugenio et al. (2023), respectively. The combination of the exposure for the eight dither positions is done with a drizzle weighting method that allows us to sub-sample the detector pixels (Law et al. 2023), resulting in cube spaxels with a size of (corresponding to pc/spaxel), and covering a Field of View (FoV) of . The absolute astrometric registration is performed using ALMA high-resolution maps of millimeter continuum emission, as described in Perna et al. (2024). The final data cubes showed strong “wiggles”, sinusoidal modulations caused by the undersampling of the point-spread function (PSF). We corrected these artifacts by applying the methodology described in Perna et al. (2023).
Figure 1 shows the three synthetic narrow-band images generated from the reduced data-cube of Arp 220 obtained for the brightest emission lines in the wavelength range covered by NIRSpec. The positions of the W and E nuclei are marked for visual purposes. This figure illustrates the quality of NIRSpec data and provides a preview of the intricate gas kinematics within the innermost regions of Arp 220.
@@ 3 Spectral fitting analysis: stellar and gas components
The analysis is divided into three steps: the first is the modelling and the subtraction of the stellar continuum, the second is a preliminary fit of the emission lines to build a model that can be used to study the integrated properties of each emission line. The third step is the disentangling of the different kinematic components associated with the most prominent emission lines. These three steps are described in detail in the following sections.
3.1 Modelling of the stellar continuum
To fit the stellar continuum, we used the penalized pixel-fitting routine (pPXF) (Cappellari & Emsellem 2004; Cappellari 2017) to convolve stellar spectra templates with a Gaussian velocity distribution. We binned spaxels together using the Voronoi binning scheme of Cappellari & Copin (2003), to achieve a minimum S/N. We established a target minimum S/N per wavelength channel for each bin (following e.g. Belfiore et al. 2019). We used the MARCS stellar population synthesis templates (Gustafsson et al. 2008) to model the stellar continuum; these templates cover the wavelength range 0.96 – 20 m with a constant spectral resolution = 20000. We also adopted a third-order multiplicative and additive polynomials to better reproduce the spectral continuum shape. To improve the quality of the fit, we included in the total model the most prominent emission lines (see Table 1) parameterized with two Gaussian components.
@@ 3.2 Multi-Gaussian emission line fit
After the subtraction of the continuum, we smoothed each wavelength map of the cube (i.e. image generated with a spectral pixel wide synthetic filter) with a Gaussian having a dispersion of 1 spaxel (i.e. 0.05″), to enhance the S/N. We then performed separate spectral fits for the atomic and molecular gas line groups listed in Table 1 since each gas phase may be dominated by different kinematic components, as illustrated in Fig. 1 (see also e.g., May & Steiner 2017; Perna et al. 2024). All lines within each group were simultaneously fitted using one, two, or three Gaussian components. We tied the velocity and the velocity dispersion of each Gaussian component among all lines of the group, while the fluxes are left free to vary. Then, we chose the number of Gaussian components for each spaxel by performing a Kolmogorov-Smirnov (KS) test on the distribution of the residuals with and components to define the minimum number of components required to provide an acceptable fit in each spaxel (Marasco et al. 2020). Consequently, the number of Gaussian components is not based on a physical model but it is constrained on just a statistical analysis.
Figure 4 shows the three moment maps of the brightest emission lines: Hm, Pa, and [Fe ii] 1.257 obtained from the fit. These maps allowed us to identify the different kinematic and morphological components of the system, traced by ionised and molecular gas, and described in Sect. 4.
@@ 3.3 Systemic velocity
In this work, the large-scale ( pc) stellar and gas kinematics maps are derived by using as a reference the redshift reported in Perna et al. (2020), which corresponds to the median value between and , (e.g. Fig. 4). This redshift allows us to obtain a symmetric stellar velocity gradient, associated with the kpc-scale disk also identified with optical and cold molecular line tracers (see Perna et al. 2020 and references therein). In contrast, the small-scale motions of the ionized and hot molecular gas around the E or W nucleus are measured relative to the zero velocity inferred from and , respectively.
@@ 4 Spectral fitting analysis: separating rotations and outflows components
The morphology and kinematics of the nuclear region of Arp 220 are quite complex. Figs. 1 and 4 reveal a superposition of circum-nuclear and large-scale disks, as well as multiple outflows already observed in optical and sub-mm tracers (e.g. Scoville et al. 2017; Perna et al. 2020; Lamperti et al. 2022; Ueda et al. 2022). Emission lines often display multi-peak and asymmetric spectral profiles, requiring the use of multiple Gaussian components for modelling. Conventional methods usually employed to separate a disk from an outflow structure across the entire galaxy (e.g. based on velocity or velocity dispersion criteria; Tozzi et al. 2021; Venturi et al. 2021), may not be straightforward for Arp 220. Hence, in this work we opted for an alternative approach, focusing on individual regions where the separation of distinct kinematic components is facilitated by dedicated criteria and by the high spatial resolution of NIRSpec data. In particular, we first selected and modelled the most easily distinguishable outflow components, subtracted them, and then gradually move on to the more complex ones.
Our method consists of three steps. The first is the identification of a region with specific morphology and/or kinematics, and hence the construction of a spatial mask to isolate it. We then fitted all spaxels within the selected region with N Gaussians, where N represents the number of the kinematically distinct components we wanted to characterise. As the fit of a complex line profile is subject to degeneracy, the value N is quite critical and depends on the detectability of the morphological and kinematic features which we aim to describe. We considered in our model the most intense transitions of the atomic (Pa) and molecular gas (Hm, hereafter H2 unless otherwise specified) in the G235H cube. We also included in the fit the HeI 1.869 m and He 1.870 m lines since they can contaminate the profile of Pa (see e.g. Fig. B.3 in Perna et al. 2024). We fitted together the transitions of atomic and molecular gas assuming for each line the same and and varying the fluxes. As we will show in the next sections, this is a reasonable assumption for this system. In the second step, we defined a criterion to isolate an outflow among the components identified in the mask. Then, we constructed a model cube containing only the isolated component, used for further analysis presented in the following sections.
@@
Fig. 5a illustrates the masks for all the identified outflow regions. They encompass a conical outflow launched by the E nucleus (i.e. the south-east outflow, ‘SEO’ hereinafter), an hourglass-shaped outflow originating from the W nucleus (‘HGO’), and a north-western bubble (‘NWB’). There is another feature of particular interest, affecting just a few spaxels and therefore not reported in the figure: a compact outflow observed in the W nucleus mainly in the ionised phase (‘WNO’). In the next subsections, we provide detailed analyses of each identified region.
4.1 South-East conical outflow from E Nucleus
Figure 5b shows H2 emission in the velocity range km s-1 with respect to the systemic velocity of the E nucleus. In this figure, a high-velocity conical region situated south-east of the E nucleus is clearly visible (demarcated by white lines). This region has also elevated velocity dispersion ( km s-1), as shown in the moment-2 maps of both ionized and hot molecular gas in Fig. 4, hence indicating the presence of an outflowing component.
We selected the region with enhanced flux in the km s-1 map (red area in Fig. 5a) and, within this region, fitted both the ionized gas and the molecular gas with three Gaussian components to model: 1) the high-velocity gas, 2) the rotating disc at a large scale that covers the entire FoV, and 3) the emission from the small circum-nuclear disc around the E nucleus. Among these three Gaussian components, we assigned the most blue-shifted to the outflow. The flux, the velocity and the velocity dispersion of the SEO in H2 and Pa are reported in Fig. 6. Since we associated only one Gaussian component to the outflow, the velocity and the velocity dispersion is the same for both ionised and hot molecular gas. The velocity varies between km s-1, observed close to the nucleus and at the edges of the cone, and km s-1 at larger distances ( pc). High velocity dispersions (up to 200 km s-1) are observed in the center of the conical outflow and close to the nucleus, while at the edges of the cone is about two times lower.
Figure 7 shows the spectra of the hot molecular (top) and ionized gas (bottom) extracted from three spaxels within the conical outflow; specifically near the E nucleus, at the center, and at the edge of the cone. In all spaxels, we observe a prominent wing extending towards negative velocities; this wing is reproduced by a Gaussian component with km s-1, which we associated with the SEO (red curves in the figure). We also note that the SEO emission is more prominent and contributes more to the total line profile in the hot molecular phase than in the ionized one, especially in the off-nuclear regions. Figure 7 also shows that our multi-Gaussian approach is capable of simultaneously reproducing the profiles of ionized and hot molecular gas, hence supporting the validity of our assumption about the presence of a mixed, multiphase ISM in each kinematic component.

### v2

@@
The study of starburst and active galactic nuclei (AGN) feedback is crucial for understanding the regulation of star formation and the evolution of galaxies across cosmic time. Arp 220, the closest ultraluminous infrared galaxy (ULIRG), is in an advanced phase of a major merger with two distinct nuclei, and it shows evidence of multiphase (molecular, ionized, and neutral) and multiscale (from ¡ 0.1 to ¿ 5 kpc) outflows. Therefore, it represents an ideal system for investigating outflow mechanisms and feedback phenomena in detail. Using new JWST NIRSpec IFU observations, we investigated the spatially resolved gaseous (in both ionized and hot molecular phases) and stellar kinematics in the innermost 1 kpc. We decoupled the different gas kinematic components through multi-Gaussian fitting, identifying two multiphase outflows, each associated with one nucleus, with velocities up to km s-1. We also resolved two counter-rotating discs around each nucleus embedded in a larger-scale rotational disk. We compute the total (including ionized, cold, and hot molecular) outflow mass ( M⊙), the mass rate ( M⊙ yr-1), and the energetics ( erg s-1) for each nucleus, and we found that the ionized and hot molecular outflowing gas contribute around 2-30% of the total mass and the energy of the outflows, as inferred from the combination of multiwavelength information. We discuss the possible origin of the outflows, finding no compelling evidence to prefer a starburst- or AGN-driven scenario. Regardless of their nature, outflows in Arp 220 propagate in multiple directions from parsec to kiloparsec scales, potentially impacting a significant portion of the host galaxy. This contrasts with isolated systems where outflows typically follow a more collimated path or are limited to the central region of the galaxy and hence do not affect the interstellar medium throughout the entire galaxy. This study highlights the importance of investigating merging systems with multiwavelength facilities, including JWST/NIRSpec IFU, to obtain a comprehensive understanding of feedback mechanisms in galaxy evolution.
@@
Ultraluminous infrared galaxies (ULIRGs) with infrared (IR) luminosities LIR ¿ L⊙ are among the most luminous objects in the local Universe. Their extreme IR luminosity is due to the combination of high star formation rate (SFR) and active galactic nuclei (AGN) activity, coupled with a high dust content, which intercepts the UV photons emitted from either hot young stars or AGN and reemits them at longer wavelengths (Soifer et al. 1984; Sanders & Mirabel 1996; Lonsdale et al. 2006; U et al. 2012 and references therein).
@@
In the starburst-quasar evolutionary scenario (e.g., Sanders et al. 1988; Hopkins et al. 2008), ULIRGs are thought to represent a rapidly growing phase of massive galaxies. Within this framework, the merger of gas-rich galaxies triggers a phase of vigorous star formation (SF), fueling the growth of supermassive black holes (SMBHs) and the eventual onset of AGN activity. As the merger progresses, powerful outflows generated by both starburst and AGN winds shape the galactic environment, influencing subsequent SF and black hole accretion with feedback mechanisms (e.g., Di Matteo et al. 2005; Silk 2013).
@@
Galactic outflows are expected to be more important at high redshift () during the peak of the cosmic SF and SMBHs accretion (Madau & Dickinson 2014), and hence where the feedback mechanisms may be maximized (e.g., Brusa et al. 2015; Talia et al. 2017; Perrotta et al. 2019; Kakkad et al. 2020; Cresci et al. 2023). Remarkably, recent observations have revealed some similarities between local ULIRGs and distant starbursts at redshift (e.g., Arribas et al. 2012; Hung et al. 2014; Bellocchi et al. 2022; McKinney et al. 2023) and they highlight the importance of understanding the outflow phenomena occurring in local ULIRGs. In fact, the proximity of these sources allows for more detailed observations that are characterized by angular resolution and signal-to-noise ratios higher than those focusing on more distant galaxies.
@@
Several studies have demonstrated the presence of multiphase (ionized, neutral, and molecular) gas outflows in ULIRGs (e.g., Sturm et al. 2011; Bellocchi et al. 2013; Veilleux et al. 2013; Arribas et al. 2014; Feruglio et al. 2015; Cazzoli et al. 2016; Fluetsch et al. 2021; Perna et al. 2021; Lai et al. 2022; Bohn et al. 2024). In this work, we focus on the closest prototypical ULIRG, Arp 220, where evidence of multiphase outflows has already been obtained from several multiwavelength observations (e.g., Arribas et al. 2001; McDowell et al. 2003; Colina et al. 2004; Perna et al. 2020; Wheeler et al. 2020; Lamperti et al. 2022; Ueda et al. 2022).
@@
Arp 220 is located at a distance of Mpc, and has an IR luminosity of log () = 12.2 (Pereira-Santaella et al. 2021), and a stellar mass of log (/ (U et al. 2012). It is a late-stage merger containing two compact ( pc), highly obscured nuclei (east and west) separated by 370 pc (1″) that dominate the IR luminosity (e.g., Scoville et al. 2007). The two nuclei are sites of intense SF, with SFR M⊙ yr-1 (e.g., Nardini et al. 2010; Varenius et al. 2016) and extremely high SFR surface densities M⊙ yr-1 kpc2, from radio and far-IR (FIR) observations (e.g., Barcos-Muñoz et al. 2015; Pereira-Santaella et al. 2021). There is still no convincing direct evidence for the presence of AGN activity in either of the two nuclei (see Perna et al. 2024 and references therein).
@@
This paper presents a kinematic examination of the nuclear region of Arp 220 observed with the Integral Field Spectrograph (IFS) unit of the Near Infrared Spectrograph (NIRSpec) instrument on board the James Webb Space Telescope (JWST) (Jakobsen et al. 2022; Böker et al. 2022) collected as part of the JWST/NIRSpec IFS Guaranteed Time Observations (GTO) survey: Resolved structure and kinematics of the nuclear regions of nearby galaxies (program lead: Torsten Böker). Our goal is to investigate the spatially resolved gaseous and stellar kinematics in the innermost 1 kpc of Arp 220, and better understand the feedback processes affecting this ULIRG system.
@@
This paper is organized as follows. Section 2 presents the NIRSpec observations and data reduction. Sections 3 and 4 present the spectral fitting analysis and the results through the characterization of the spatially resolved kinematics of the atomic and molecular gas and the identification and study of the different structures found in the central 1.5 1.8 kpc of Arp 220. In Sect. 5, we describe the main properties of the interstellar medium (ISM). In Sect. 6, we discuss our results and a comparison with the literature. Finally, Sect. 7 summarizes our conclusions. Throughout this work, we assume and km s-1 Mpc-1 (Bennett et al., 2014). With this cosmology, corresponds to 0.368 kpc at the distance of Arp 220.
@@ 2 Observation and data reduction
Arp 220 was observed using the JWST/NIRSpec instrument in IFS mode (Böker et al. 2022), as part of the JWST observing program 1267 (PI: D. Dicken), on March 6, 2023. These NIRSpec observations were combined in a single proposal with independent MIRI GTO observations of the same target.
The dataset comprises two distinct pointings, each employing the first four positions of the medium cycling dithering pattern with 15 groups per integration and one integration per exposure, using the NRSIRS2RAPID readout pattern. The first set is centered at the midpoint between the coordinates of the western (W) and eastern (E) nuclei, while the second set is offset by approximately 1″ toward the northwest to encompass the H and [N ii] emission shell (Lockhart et al. 2015; Perna et al. 2020). Each set has a total integration time of 933 seconds and utilises three high-resolution (R ) configurations: G140H/F100LP, G235H/F170LP, and G395H/F290LP, covering wavelength ranges of 0.97 – 1.82 m, 1.66 – 3.05 m, and 2.87 – 5.14 m, respectively. However, this paper focuses solely on the first two configurations, as our objective is to study the kinematics of the emitting gas, with its strongest transitions occurring at wavelengths below 2.4 m.
The data reduction of Arp 220 observations is described in detail in Perna et al. (2024). Here we summarize the relevant information. We used the JWST pipeline v1.8.2 with CRDS context 1063. A patch was included to correct some important bugs that affect this specific version of the pipeline; noise and outliers were corrected following the approaches described in Perna et al. (2023) and D’Eugenio et al. (2023), respectively. The combination of the exposure for the eight dither positions was done with a drizzle weighting method that allowed us to sub-sample the detector pixels (Law et al. 2023), resulting in cube spaxels with a size of (corresponding to pc/spaxel) and covering a field of view (FoV) of . The absolute astrometric registration was performed using ALMA high-resolution maps of millimeter continuum emission, as described in Perna et al. (2024). The final data cubes showed strong “wiggles”, sinusoidal modulations caused by the undersampling of the point spread function (PSF). We corrected these artifacts by applying the methodology described in Perna et al. (2023).
Figure 1 shows the three synthetic narrow-band images generated from the reduced data cube of Arp 220 obtained for the brightest emission lines in the wavelength range covered by NIRSpec. The positions of the W and E nuclei are marked for visual purposes. This figure illustrates the quality of NIRSpec data and provides a preview of the intricate gas kinematics within the innermost regions of Arp 220.
@@ 3 Spectral fitting analysis: Stellar and gas components
The analysis is divided into three steps: the first is the modeling and the subtraction of the stellar continuum, the second is a preliminary fit of the emission lines to build a model that can be used to study the integrated properties of each emission line. The third step is the disentangling of the different kinematic components associated with the most prominent emission lines. These three steps are described in detail in the following sections.
3.1 Modeling of the stellar continuum
To fit the stellar continuum, we used the penalized pixel-fitting routine (pPXF) (Cappellari & Emsellem 2004; Cappellari 2017) to convolve stellar spectra templates with a Gaussian velocity distribution. We did not attempt to extract the Gauss–Hermite terms h3 and h4, since the quality of our data is not sufficient to obtain this information from the fit (see also e.g., Cappellari et al. 2009; Engel et al. 2011a; Crespo Gómez et al. 2021). We binned spaxels together using the Voronoi binning scheme of Cappellari & Copin (2003), to achieve a minimum S/N. We established a target minimum S/N per wavelength channel for each bin (following e.g., Belfiore et al. 2019). We used the MARCS stellar population synthesis templates (Gustafsson et al. 2008) to model the stellar continuum; these templates cover the wavelength range 0.96 – 20 m with a constant spectral resolution = 20000. We also adopted a third-order multiplicative and additive polynomials to better reproduce the spectral continuum shape. To improve the quality of the fit, we included in the total model the most prominent emission lines (see Table 1) parameterized with two Gaussian components.
@@ 3.2 Multi-Gaussian emission line fit
After the subtraction of the continuum, we smoothed each wavelength map of the cube (i.e., image generated with a spectral pixel wide synthetic filter) with a Gaussian having a dispersion of 1 spaxel (i.e., 0.05″), to enhance the S/N. We then performed separate spectral fits for the atomic and molecular gas line groups listed in Table 1 since each gas phase may be dominated by different kinematic components, as illustrated in Fig. 1 (see also e.g., May & Steiner 2017; Perna et al. 2024). All lines within each group were simultaneously fitted using one, two, or three Gaussian components. We tied the velocity and the velocity dispersion of each Gaussian component among all lines of the group, while the fluxes are left free to vary. Then, we chose the number of Gaussian components for each spaxel by performing a Kolmogorov-Smirnov (KS) test on the distribution of the residuals with and components to define the minimum number of components required to provide an acceptable fit in each spaxel (Marasco et al. 2020). Consequently, the number of Gaussian components is not based on a physical model but it is constrained on just a statistical analysis; a detailed, physical decomposition is presented in Sect. 4.
Figure 4 shows the three moment maps calculated on the whole modeled line profile (consisting of either one, two, or three Gaussians) of the brightest emission lines: Hm, Pa, and [Fe ii] 1.257. For each transition, we masked spaxels with S/N3. These maps allowed us to identify the different kinematic and morphological components of the system, traced by ionized and molecular gas, and described in Sect. 4.
@@ 3.3 Systemic velocity
In this work, the large-scale ( pc) stellar and gas kinematics maps are derived by using as a reference the redshift reported in Perna et al. (2020), which corresponds to the median value between and , (e.g., Fig. 4). This redshift allows us to obtain a symmetric stellar velocity gradient, associated with the kiloparsec-scale disk also identified with optical and cold molecular line tracers (see Perna et al. 2020 and references therein). In contrast, the small-scale motions of the ionized and hot molecular gas around the E or W nucleus are measured relative to the zero velocity inferred from and , respectively.
@@ 4 Spectral fitting analysis: Separating rotations and outflows components
The morphology and kinematics of the nuclear region of Arp 220 are quite complex. Figs. 1 and 4 reveal a superposition of circum-nuclear and large-scale disks as well as multiple outflows that have already been observed in optical and submillimeter tracers (e.g., Scoville et al. 2017; Perna et al. 2020; Lamperti et al. 2022; Ueda et al. 2022). Emission lines often display multipeak and asymmetric spectral profiles, requiring the use of multiple Gaussian components for modeling. Conventional methods usually employed to separate a disk from an outflow structure across the entire galaxy (e.g., based on velocity or velocity dispersion criteria; Tozzi et al. 2021; Venturi et al. 2021), may not be straightforward for Arp 220. Hence, in this work we opted for an alternative approach, focusing on individual regions where the separation of distinct kinematic components is facilitated by dedicated criteria and by the high spatial resolution of NIRSpec data. In particular, we first selected and modeled the most easily distinguishable outflow components, subtracted them, and then gradually move on to the more complex ones.
Our method consists of three steps. The first is the identification of a region with specific morphology and/or kinematics, and hence the construction of a spatial mask to isolate it. We then fitted all spaxels within the selected region with N Gaussians, where N represents the number of the kinematically distinct components we wanted to characterize. As the fit of a complex line profile is subject to degeneracy, the value N is quite critical and depends on the detectability of the morphological and kinematic features we aim to describe. We considered in our model the most intense transitions of the atomic (Pa) and molecular gas (Hm, hereafter H2 unless otherwise specified) in the G235H cube. We also included in the fit the HeI 1.869 m and He 1.870 m lines since they can contaminate the profile of Pa (see, e.g., Fig. B.3 in Perna et al. 2024). We fitted together the transitions of atomic and molecular gas assuming for each line the same and and varying the fluxes. As we show in the next sections, this is a reasonable assumption for this system. In the second step, we defined a criterion to isolate an outflow among the components identified in the mask. Then, we constructed a model cube containing only the isolated component, used for further analysis presented in the following sections.
@@
Fig. 5a illustrates the masks for all the identified outflow regions. They encompass a conical outflow launched by the E nucleus (i.e., the southeast outflow, SEO hereinafter), an hourglass-shaped outflow originating from the W nucleus (HGO), and a northwestern bubble (NWB). There is another feature of particular interest that affects just a few spaxels and is therefore not reported in the figure: a compact outflow observed in the W nucleus mainly in the ionized phase (WNO). In the next subsections, we provide detailed analyses of each identified region.
4.1 Southeast conical outflow from E nucleus
Figure 5b shows H2 emission in the velocity range km s-1 with respect to the systemic velocity of the E nucleus. In this figure, a high-velocity conical region situated southeast of the E nucleus is clearly visible (demarcated by white lines). This region has also elevated velocity dispersion ( km s-1), as shown in the moment-2 maps of both ionized and hot molecular gas in Fig. 4, hence indicating the presence of an outflowing component.
We selected the region with enhanced flux in the km s-1 map (red area in Fig. 5a) and, within this region, fitted both the ionized gas and the molecular gas with three Gaussian components to model: 1) the high-velocity gas, 2) the rotating disk at a large scale that covers the entire FoV, and 3) the emission from the small circum-nuclear disk around the E nucleus. Among these three Gaussian components, we assigned the most blueshifted to the outflow. The flux, the velocity and the velocity dispersion of the SEO in H2 and Pa are reported in Fig. 6. Since we associated only one Gaussian component to the outflow, the velocity and the velocity dispersion is the same for both ionized and hot molecular gas. The velocity varies between km s-1, observed close to the nucleus and at the edges of the cone, and km s-1 at larger distances ( pc). High velocity dispersions (up to 200 km s-1) are observed in the center of the conical outflow and close to the nucleus, while at the edges of the cone is about two times lower.
Figure 7 shows the spectra of the hot molecular (top) and ionized gas (bottom) extracted from three spaxels within the conical outflow; specifically near the E nucleus, at the center, and at the edge of the cone. In all spaxels, we observe a prominent wing extending toward negative velocities; this wing is reproduced by a Gaussian component with km s-1, which we associated with the SEO (red curves in the figure). We also note that the SEO emission is more prominent and contributes more to the total line profile in the hot molecular phase than in the ionized one, especially in the off-nuclear regions. Figure 7 also shows that our multi-Gaussian approach is capable of simultaneously reproducing the profiles of ionized and hot molecular gas, hence supporting the validity of our assumption about the presence of a mixed, multiphase ISM in each kinematic component.